OpenStudy (mysesshou):

Find the limit of nth sqrt(n) So, I was just wondering which method is best and if there is any better way of doing this. Thanks guys.

OpenStudy (eyust707):

okay so im not sure of the techincal way of doing the first one but intuitively if we look at the graph of ln (x) it looks like this

OpenStudy (eyust707):



OpenStudy (eyust707):

and yes as x goes off to infinity, so does y, but not very fast

OpenStudy (eyust707):

infinity therefore gets bigger faster than ln(infinty) does

OpenStudy (eyust707):

so you end up with a smaller number over a bigger number which goes off to zero..

OpenStudy (eyust707):

hehe i think...

OpenStudy (mysesshou):

Thx. Thinking in graph method is more easily comprehended sometimes. :)

OpenStudy (eyust707):

looks like your better than i at these but once again for the second one i would just use intutition... as you said: the root n of n = n^(1/n) , as n goes to infinity th root becomes 0. anything to the zero power is 1

OpenStudy (mysesshou):

Thanks. :)

OpenStudy (eyust707):

np

OpenStudy (turingtest):

the way these should be done are on the attachments at the top of the thread the first problem is done using l'Hospitals rule the second is done by rewriting it as\[\huge \lim_{n\to\infty}\sqrt[n]n=e^{\lim_{n\to\infty}\ln \sqrt[n]n}\]we can then use log properties, then l'hospital

OpenStudy (mysesshou):

I don't think I follow why the limit can just jump up on the e. :(

OpenStudy (turingtest):

because we have that\[\huge a^{\log_a x}=x\]the trick is to go\[\huge \lim_{n\to a}f(x)=\lim_{n\to a}e^{\ln f(x)}\]since e is a constant we can just move the limit to the exponent\[\huge=e^{\lim_{n\to a}\ln f(x)}\]then hopefully you can evaluate the exponent it is okay to do this because

OpenStudy (turingtest):

strike the "it is okay to do this because" at the bottom there...

OpenStudy (anonymous):

the "e thing" is because the precise definition of \[b^x\] is \[e^{x\ln(b)}\] and you can take the limit up into the exponent because \[e^x\] is a continuous function, meaning \[\lim_{x\to a}e^{f(x)}=e^{\lim_{x\to a}f(x)}\]

OpenStudy (anonymous):

what turning test said

OpenStudy (anonymous):

well, except for the "e is a constant" part. limit goes up because \[\exp\] is a continuous function

OpenStudy (mysesshou):

Ah, right. I'm just slow and blind. I missed the ln in the exponent. I should have remembered the ln and e rules better. The last time I used them I had to look them up. Thanks so much. You're the best !!

OpenStudy (turingtest):

Thanks for correcting my reasoning @satellite73 , I shouldn't have said that since I wasn't sure.

OpenStudy (mysesshou):

Thanks you guys. You're both so helpful. :) I think I remember the log and ln rules more. Guess I should review some more.